OpenStudy (anonymous):

Solve: x+sqrt(x) = 20

OpenStudy (anonymous):

\[x + \sqrt{x}\]

OpenStudy (anonymous):

Could you show us what you've done with the question please?

OpenStudy (anonymous):

= 20

OpenStudy (anonymous):

i haven't done anything i don't know how to start it lol sorry

OpenStudy (anonymous):

Well firstly, have you learnt substitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, well we should let \[u=\sqrt{x}\]

OpenStudy (anonymous):

So everytime you see \sqrt{x} or x(which is ideally (\sqrt{x})^2 substitute u in.

OpenStudy (anonymous):

ok so then u^2 + u = 20

OpenStudy (anonymous):

Yep then move the 20 to the LHS and solve the quadratic as per usual.

OpenStudy (anonymous):

that doesn't give me the right answer

OpenStudy (anonymous):

And the after you get the two values of "u", revert back to \sqrt{x} to find the values of x.

OpenStudy (anonymous):

And then*

OpenStudy (anonymous):

that gave me 4,-5 the answers are 0 and 16

OpenStudy (anonymous):

how do i revert those back?

OpenStudy (anonymous):

We're not done yet. You havem t checked your answers back to the original equation.

OpenStudy (anonymous):

Haven't*

OpenStudy (anonymous):

Firstly. \[u=4\] or \[u=-5\]

OpenStudy (anonymous):

correct then what from there

OpenStudy (anonymous):

Then rewrite u as \sqrt{x}. So you have this: \[\sqrt{x}=4\] or \[\sqrt{x}=-5\]

OpenStudy (anonymous):

Now you square everything and you can find the two values of x.

OpenStudy (anonymous):

do you only fill it in for the sqrt root x or do you do it also for the other x that isn't sqqred

OpenStudy (anonymous):

like what should it look like?

OpenStudy (anonymous):

No, let's do the first one. What's x for the first one. \[\sqrt{x}=4\] \[x=?\]

OpenStudy (anonymous):

i have this right now